An empty railroad car of mass 4400 kg is being pulled beneath a grain elevator that dumps grain at the rate of 175 kg/s. If the coefficient of friction between the car wheels and the track is 0.021, what is the force needed to keep the car moving at a constant speed of 1.9 m/s 25.0 s after the grain is first dumped on the car

Points given for correct answer and explanation
Answer #1

Mass of Car=4400 kg

Mass of grain of car=175*25=4375 kg

Total mass=4400+4375=8775 kg

Friction force acting=mu*m*g

mu=0.021

Force required=Force of friction=0.021*8775*9.8

                                                =1806 N